OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24.

OpenStudy (lgbasallote):

hint: y varies directly with x means \[y = kx\] does that help?

OpenStudy (anonymous):

no

hartnn (hartnn):

find k from y=kx by putting y=12,x=16 and tell what u got as k.....

OpenStudy (anonymous):

ok.........one sec

OpenStudy (anonymous):

12 = k16?

hartnn (hartnn):

yes,so how do u get k on one side only??

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

divide by 12?

OpenStudy (ghazi):

Y=k*X...k=Y/X \[k=\frac{ 12 }{ 16 }\] now find K and put it in \[Y=k*X\] you've K and you have X..find Y

hartnn (hartnn):

since u want to isolate k , u divide by the number which is with k.... which number is with k?

OpenStudy (anonymous):

16

hartnn (hartnn):

yes,so u divide by 16 on both sides....

OpenStudy (anonymous):

1.3........?

OpenStudy (ghazi):

no

hartnn (hartnn):

how did u get 1.3?

OpenStudy (ghazi):

try to figure out again

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

so whats k?

OpenStudy (anonymous):

32